OpenStudy (anonymous):

r=2-6 cos(θ) convert this polar coordinate equation into an equation in rectangular coordinates(x,y) given (-8,8) convert this point into polar coordinates with 6 >= 0 and 0<=θ<= 2pi given x^2+ (y-5)^2=25 cover this rectangular coordinate equation in to an equation in polar coordinates (r,θ)

myininaya (myininaya):

cos(theta)=x/r sin(theta)=y/r r^2=x^2+y^2 ----------------------------- so we have r=2-6cos(theta) sqrt{x^2+y^2}=2-6*x/r

myininaya (myininaya):

but r=sqrt{x^2+y^2} sqrt{x^2+y^2}=2-6*x/(sqrt{x^2+y^2}

myininaya (myininaya):

if you want you can multiply both sides of that equation by sqrt{x^2+y^2}

myininaya (myininaya):

for the second one we have (x,y)=(rcostheta,rsintheta) rcostheta=x rsintheta=y costheta=x/r sintheta=y/r draw a traingle in the second quadrant you should get the hypethuese or r=8sqrt{2} and you should get theta=3pi/4

myininaya (myininaya):

for the last one just use rcostheta=x and rsintheta=y r^2cos^2theta+(rsintheta-5)^2=25 now multiply (rsintheta-5)(rsintheta-5) and you should be able to use cos^2x+sin^2x=1 eventually